OpenStudy (xx_gamergirl_xx):

FAN & MEDAL FOR HELP PLZ HELP I HAVE TO TURN THIS IN TODAY -_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_- Find both unit rates. 1435 points scored in 25 games A. 57.4 points/game and about 0.017 game/points B. 57.4 game/points and about 0.017 point/game C. 1435 points/game and 25 games/point D. 1435 games/point and 25 points/game I Have No Idea What The Answer Could Be

OpenStudy (superdavesuper):

Almost exactly the same type of question as last one....plz take a look from the last solution

OpenStudy (xx_gamergirl_xx):

ik but still I did but I still am confused soz if im confused and im being rushed and I cant do good when im being rushed

OpenStudy (superdavesuper):

1435/25 is one of the answers...

OpenStudy (xx_gamergirl_xx):

thx I guess

OpenStudy (xx_gamergirl_xx):

@Tootles143 can u help plz im confused with this

OpenStudy (xx_gamergirl_xx):

so do 1435/25 I did and got 457.4

OpenStudy (xx_gamergirl_xx):

@inowalst can u help plz im sooo confused still

OpenStudy (inowalst):

@TheSmartOne Would you mind again?

OpenStudy (inowalst):

I dont get this.

OpenStudy (xx_gamergirl_xx):

me ethier and its my math :(

OpenStudy (xx_gamergirl_xx):

@iambatman can u help plz

OpenStudy (xx_gamergirl_xx):

@Hero can u help plz

OpenStudy (xx_gamergirl_xx):

@wio can u help plz

TheSmartOne (thesmartone):

1435 points scored in 25 games So to get how many points he/she get in one game then we are looking for: x points scored in 1 game So we can make a ratio: \(\sf\Large\frac{1435}{25}=\frac{x}{1}\)

OpenStudy (xx_gamergirl_xx):

ohhhh

OpenStudy (xx_gamergirl_xx):

so about 57 points in 1 game?

TheSmartOne (thesmartone):

And solve for x. And you get the first half of your answer.

OpenStudy (xx_gamergirl_xx):

so 57/1 ?

OpenStudy (xx_gamergirl_xx):

er 57.4/1

OpenStudy (xx_gamergirl_xx):

ohh

TheSmartOne (thesmartone):

.And then to find how many games per point. We have 1435 points scored in 25 games and we want 1 point per y games We get the proprtion of \(\sf\Large\frac{1435}{25}=\frac{1}{y}\) And then we just cross multiply to find y :)

OpenStudy (xx_gamergirl_xx):

oh ok so then it would be A?

TheSmartOne (thesmartone):

And you answer will be in the form of: x points/game, y game/points

OpenStudy (xx_gamergirl_xx):

I mean B